New ZOOM room:

News:

shift at P5 : Patrizia should re-vamp her access requests Lucia pushes for remote shifters : DAQ and DQM

Framework & production

Federica : /lsf will not be longer accessible after end of March. Condor will be the supported alternative. Instruction will be

HTCondor-howto (in progress) direct link it should be added to the CMS Padova webpage

Analysis

B physics

Enrico : J/PsiPhi GL and it will be submitted soon wink Hevjin : presentation about her status in one of the upcoming Padova CMS meeting --> contact Roberto

Heavy composite neutrinos

Patrizia : Matteo submitted his final version of the thesis. Unblinded data plots will be presented soon electron channel is ok, muon channel has probably a small bug (which has been highlighted by the impact of the systematics in the combination). Next week a presentation for switching the pre-approval flag.

VBS

Patrizia : on Tue presentation on the Z channel. the current status suggests the analysis needs to move to UL. Matteo will have "assegno di ricerca" in Perugia.

WH_SS

Siew Yan : 3L channel has some excess which needs to be investigated (ongoing and followed by the 3L channel team). The same-sign is under control, instead. For the ZH channel converging on the answering to the comments received during approval.

H/Z->JPsi gamma

Alberto : signal w/ madgraph has been approved. Gaia is working on the photon regression; Alberto is studying the categorisation approach for improving the signal significance

EPR

Anushree : ML application to the data certification presented at the tracker DQM meeting

Enrico : miniAOD

Detector

Trigger

Jacopo : evaluate performance of the track reconstruction. Issues in the communication of the vado boards, Jacopo will .

Timing Layer

Roberto : Valentina Sola presented MTD at Gr1, in this occasion the TEC has been presented officially to the INFN management as well (the TEC has been already approved by CMS) 200k CHF are requested to INFN and we will see after the "private" meeting next week. The fridge is in the workshop, Roberto Isocrate will bring the rack to the lab, so possibly next week we will have the whole components for starting the measurement of the cooling plate. Confirmation from the company about the installation of the TEC : 2/3 weeks they will be back.

DT

Fabrizio : AN about the ageing studies performed in the last years in Legnaro, it will be useful for the TIF studies. Version this week.

Sandro : focusing on upgrade. the workshop about integration and mechanics will continue next week

Tracker/tracking

ntr

A.O.B.

-- MiaTosi - 2021-02-16

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2021-02-16 - MiaTosi
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback